OpenStudy (anonymous):

if y=5 find the value of 25/y

OpenStudy (mimi_x3):

5 xD

OpenStudy (katrinakaif):

5

OpenStudy (mathhelp346):

5

jimthompson5910 (jim_thompson5910):

25/y is the same as 25/5 because y is 5 So 25/y = 25/5 = 5

OpenStudy (saifoo.khan):

\[\Huge \frac{25}{5}=\frac{\cancel{25}}{\cancel{5}}=5\]